# Tree Matching Networks for Natural Language Inference: Parameter-Efficient Semantic Understanding via Dependency Parse Trees

### v1

@@ Abstract
However, we identify a scaling plateau: increasing TMN parameters 600 yields minimal improvement, indicating that simple aggregation methods create an architectural bottleneck. These findings motivate hybrid approaches that preserve structural benefits while addressing scalability limitations, such as using graph neural networks to pre-encode dependency trees before transformer-based aggregation with tree-aware positional encodings.
@@ Introduction
TreeMatchingNet achieves 60.70% SNLI test accuracy compared to BertMatchingNet’s 35.38%, demonstrating 1.7 superior performance with comparable parameter counts. This translates to improved parameter efficiency: TMN achieves 1.96 better efficiency (1.69% vs 0.86% accuracy per million parameters). However, we observe a scaling plateau: increasing TMN parameters from 60K to 36M (600) yields minimal improvement, revealing an architectural bottleneck in the aggregation mechanism. Meanwhile, BertMatchingNet exhibits a failure mode where it predicts every test example as entailment, suggesting incompatibility between the cross-attention architecture and BERT’s design.
@@
We present an adaptation of Graph Matching Networks to linguistic dependency trees with rich features combining BERT embeddings, POS tags, and morphological annotations for natural language inference. Our evaluation compares tree-based GNN and BERT architectures at matched parameter counts (36M) with identical training protocols, enabling controlled analysis of structural inductive bias. We identify a scaling plateau where 600 parameter increases yield minimal improvement, suggesting the aggregation mechanism as an architectural bottleneck. Through systematic comparison of matching versus embedding architectures, we isolate the contribution of cross-graph attention and observe compatibility differences between tree-based and transformer-based approaches.
@@
| SNLI | Entailment | 550,152 | 10K / 10K | 540,803 | 9,969 / 9,960 | Complete dataset |
| SemEval | Similarity | 3,000 | 750 / 6,750 | 2,967 | 750 / 6,638 | |
Data Preparation. Raw text datasets were converted to dependency tree representations using TMN_DataGen (Section 3.1.1). For large-scale contrastive pretraining (WikiQS + AmazonQA), computational constraints limited full-dataset conversion and training: we converted 2.52M training pairs and randomly sample 600 batches per epoch (153,600 pairs, batch size 256) without replacement within each epoch. This sampling strategy processes 6% of the contrastive corpus per epoch. Each epoch uses a different random sample from the full pool. For task-specific datasets (SNLI, SemEval), we use complete converted datasets without sampling.
@@ 1. TreeEncoder: Maps input features to initial node and edge representations:
| TreeMatchingNet Small (60K params) | |
@@
TreeMatchingNet achieves 60.70% test accuracy compared to BertMatchingNet’s 35.38%, representing a 1.7 performance advantage with comparable parameter counts (36M vs 41M).While both models fall short of state-of-the-art performance (RoBERTa-large achieves approximately 90% with 355M parameters and extensive pretraining on 160GB of text), the comparison reveals that structural inductive biases provide substantial advantages at moderate parameter scales when training data is limited.
@@
| C | 2009 | 654 | 564 | 3227 |
| N | 1017 | 1397 | 795 | 3209 |
| E | 400 | 484 | 2640 | 3524 |
Per-class metrics (Table 6) reveal varying performance across classes. The model achieves strongest performance on Entailment (74.91% recall, 66.02% precision), adequate performance on Contradiction (62.26% recall, 58.64% precision), and weaker performance on Neutral (43.53% recall, 55.11% precision).
@@
| Contradiction | 58.64% | 62.26% | 60.39% | 3227 |
| Neutral | 55.11% | 43.53% | 48.64% | 3209 |
| Entailment | 66.02% | 74.91% | 70.18% | 3524 |
The Neutral class shows the weakest performance, with 31.7% of neutral examples misclassified as Contradiction and 24.8% misclassified as Entailment. This pattern is consistent with the inherent ambiguity of the Neutral category, which is defined by the absence of a clear relationship rather than the presence of one.
@@
Table 10 shows that increasing TreeMatchingNet parameters from 60K to 36M (a 600 increase) yields minimal performance improvement.
@@
| TMN (Small) | 60K | 60% |
| TMN (Large) | 36M | 60.70% |
This scaling plateau indicates an architectural bottleneck preventing effective utilization of additional parameters. In typical deep learning scenarios, increasing parameters by 600 with sufficient training data yields substantial improvements. The near-zero improvement here suggests that additional capacity cannot be effectively utilized by the current architecture.
@@
| TreeEmbedding | 36M | 51% |
| BertEmbedding | 41M | 40% |
Table 9 presents preliminary results from validation curves. TreeEmbeddingNet achieves approximately 51% accuracy while BertEmbeddingNet reaches approximately 40%. The performance gap between tree-based and BERT-based models remains substantial (11 percentage points) but is noticeably smaller than the matching model gap (25.3 percentage points).
TreeEmbeddingNet outperforms BertEmbeddingNet by 19 points, confirming that the structural inductive bias of dependency trees provides benefits independent of cross-attention mechanisms. The 19-point gap for embedding models is smaller than the 25-point gap for matching models, suggesting that cross-attention amplifies the advantage of structural representations—trees benefit more from cross-graph matching than sequences do.
BertEmbeddingNet (40%) substantially outperforms BertMatchingNet (35.38%), indicating that the cross-attention mechanism interfered with BERT’s ability to learn meaningful representations. This aligns with BERT’s original design for embedding-style processing (masked language modeling, sentence pair classification without explicit cross-attention). TreeMatchingNet (60.70%) shows preforms better than TreeEmbeddingNet (51%), while BertMatchingNet shows severe degradation from BertEmbeddingNet. This asymmetry suggests fundamental architectural compatibility differences.
@@
| TreeMatching (toy) | 60K | 60% |
| TreeMatching (full) | 36M | 60.70% |
@@
TreeMatchingNet exhibits approximately the same performance between the toy model (60K parameters) and the full-sized model (36M parameters), both achieving 60% accuracy despite a 600 increase in parameters. This plateau suggests an architectural bottleneck in the aggregation mechanism that prevents effective parameter utilization, regardless of the capacity added through additional propagation layers or hidden dimensions.
@@
The embedding model results illuminate why this was so challenging for BERT: BertEmbeddingNet achieves 40% without randomized pairing, while BertMatchingNet achieves only 35.38% with randomized pairing. This 5-point degradation (13% relative) indicates that BERT’s architecture cannot effectively learn from randomized pairs processed through cross-attention. The transformer’s attention mechanisms, optimized for learning from sequential structure, fail when forced to learn from mismatched pairs where the cross-attention target is deliberately chosen to be negative.
In contrast, TreeMatchingNet (60.70%) shows no degredation from TreeEmbeddingNet (51%), instead outperforming it, indicating that graph-based message passing provides robustness to pairing randomization that transformers lack.
@@
The observed scaling plateau (60K params 36M params 60% accuracy) points to the aggregation mechanism as the primary source of information loss. The current pooling aggregation collapses node representations (each 1536-dimensional) into a single 2048-dimensional graph embedding through gated weighted sum pooling. This lossy compression represents the greatest potential for representation collapse in the architecture. While the GNN propagation layers learn rich, node-level distinctions about syntactic relationships and semantic content, the aggregation step treats all nodes similarly through a simple pooling operation, potentially destroying the structural information encoded during propagation.
@@
Both TMN models exhibit weaker performance on the Neutral class (43.53% recall) compared to Contradiction (62.26%) and Entailment (74.91%). This pattern reflects the inherent difficulty of the Neutral category.
@@ Conclusion
Key findings: (1) TreeMatchingNet achieves 60.70% SNLI accuracy vs BertMatchingNet’s 35.38%, demonstrating a clear performance advantage; (2) embedding model experiments show the structural advantage persists (TreeEmbedding 51% vs BertEmbedding 40%), with narrower gap indicating cross-attention amplifies structural benefits; (3) BertEmbeddingNet additionally outperforms BertMatchingNet, confirming that cross-attention interferes with BERT’s architecture; (4) increasing TMN parameters 600 (60K to 36M) yields minimal improvement, revealing a scaling bottleneck consistent with prior literature on strucure-based NLP approaches, but which may be overcome by introducing scalable attention-based architectures into the aggregation step of TMN.
Implications: Explicit structural representations significantly outperform sequence-based transformers at moderate scales, with benefits persisting across both matching and embedding paradigms. Inductive biases aligned with linguistic structure provide substantial learning efficiency gains. Cross-attention mechanisms amplify structural advantages for tree-based models but interfere with transformer architectures, particularly under randomized pairing regimes. Simple pooling aggregation creates bottlenecks preventing effective parameter scaling; attention-based aggregation may address this limitation.

### v2

@@ Abstract
However, we identify a scaling plateau: increasing TMN parameters 2 yields minimal improvement, indicating that simple aggregation methods create an architectural bottleneck. These findings motivate hybrid approaches that preserve structural benefits while addressing scalability limitations, such as using graph neural networks to pre-encode dependency trees before transformer-based aggregation with tree-aware positional encodings.
@@ Introduction
TreeMatchingNet achieves 75.20% SNLI test accuracy compared to BertMatchingNet’s 35.38%, demonstrating 2.1 superior performance with comparable parameter counts. This translates to improved parameter efficiency: TMN achieves 2.4 better efficiency (2.09% vs 0.86% accuracy per million parameters). However, we observe a scaling plateau: increasing TMN parameters from 18.8M to 36M (2) yields minimal improvement (0.2 percentage points), revealing an architectural bottleneck in the aggregation mechanism. Meanwhile, BertMatchingNet exhibits a failure mode where it predicts every test example as entailment, suggesting incompatibility between the cross-attention architecture and BERT’s design.
@@
We present an adaptation of Graph Matching Networks to linguistic dependency trees with rich features combining BERT embeddings, POS tags, and morphological annotations for natural language inference. Our evaluation compares tree-based GNN and BERT architectures at matched parameter counts (36M) with identical training protocols, enabling controlled analysis of structural inductive bias. We identify a scaling plateau where 2 parameter increases yield minimal improvement, suggesting the aggregation mechanism as an architectural bottleneck. Through systematic comparison of matching versus embedding architectures, we isolate the contribution of cross-graph attention and observe compatibility differences between tree-based and transformer-based approaches.
@@
| SNLI | Entailment | 550,152 | 10K / 10K | 540,803 | 9,969 / 9,960 | 600 batches/epoch (153,600 samples) |
| SemEval | Similarity | 3,000 | 750 / 6,750 | 2,967 | 750 / 6,638 | Complete dataset |
Data Preparation. Raw text datasets were converted to dependency tree representations using TMN_DataGen (Section 3.1.1). For large-scale contrastive pretraining (WikiQS + AmazonQA), computational constraints limited full-dataset conversion and training: we converted 2.52M training pairs and randomly sample 600 batches per epoch (153,600 pairs, batch size 256) without replacement within each epoch. This sampling strategy processes 6% of the contrastive corpus per epoch. Each epoch uses a different random sample from the full pool. For SNLI, we apply the same sampling strategy (600 batches per epoch) to enable comparable training across all phases. For SemEval, we use the complete dataset without sampling due to its smaller size (2,967 training pairs).
@@ 1. TreeEncoder: Maps input features to initial node and edge representations:
| TreeMatchingNet Medium (18.8M params) | |
@@
TreeMatchingNet achieves 75.20% test accuracy compared to BertMatchingNet’s 35.38%, representing a 2.1 performance advantage with comparable parameter counts (36M vs 41M).While both models fall short of state-of-the-art performance (RoBERTa-large achieves approximately 90% with 355M parameters and extensive pretraining on 160GB of text), the comparison reveals that structural inductive biases provide substantial advantages at moderate parameter scales when training data is limited.
@@
| C | 2620 | 502 | 105 | 3227 |
| N | 625 | 2086 | 498 | 3209 |
| E | 136 | 604 | 2784 | 3524 |
Per-class metrics (Table 6) reveal strong balanced performance across classes. The model achieves excellent performance on Entailment (79.00% recall, 82.20% precision), strong performance on Contradiction (81.19% recall, 77.49% precision), and good performance on Neutral (65.00% recall, 65.35% precision).
@@
| Contradiction | 77.49% | 81.19% | 79.30% | 3227 |
| Neutral | 65.35% | 65.00% | 65.18% | 3209 |
| Entailment | 82.20% | 79.00% | 80.57% | 3524 |
The Neutral class shows the weakest performance, with 19.5% of neutral examples misclassified as Contradiction and 15.5% misclassified as Entailment. This pattern is consistent with the inherent ambiguity of the Neutral category, which is defined by the absence of a clear relationship rather than the presence of one.
@@
Table 10 shows that increasing TreeMatchingNet parameters from 18.8M to 36M (approximately 2 increase) yields minimal performance improvement.
@@
| TMN (Medium) | 18.8M | 75% |
| TMN (Large) | 36M | 75.20% |
This scaling plateau indicates an architectural bottleneck preventing effective utilization of additional parameters. In typical deep learning scenarios, increasing parameters by 2 with sufficient training data yields measurable improvements. The lack of improvement here suggests that additional capacity cannot be effectively utilized by the current architecture.
@@
| TreeEmbedding | 36M | 57.57% |
| BertEmbedding | 41M | 45.78% |
Table 9 presents the final evaluation results. TreeEmbeddingNet achieves 57.57% accuracy while BertEmbeddingNet reaches 45.78%. The performance gap between tree-based and BERT-based models is 11.79 percentage points, smaller than the matching model gap of 39.82 percentage points.
TreeEmbeddingNet outperforms BertEmbeddingNet by 11.79 points, confirming that the structural inductive bias of dependency trees provides benefits independent of cross-attention mechanisms. The 11.79-point gap for embedding models is smaller than the 39.82-point gap for matching models, suggesting that cross-attention amplifies the advantage of structural representations for tree-based models more than for sequence-based models.
BertEmbeddingNet (45.78%) outperforms BertMatchingNet (35.38%) by 10.4 points, indicating that the cross-attention mechanism interfered with BERT’s ability to learn meaningful representations. This aligns with BERT’s original design for embedding-style processing (masked language modeling, sentence pair classification without explicit cross-attention). TreeMatchingNet (75.20%) outperforms TreeEmbeddingNet (57.57%) by 17.63 points, while BertMatchingNet shows degradation from BertEmbeddingNet. This asymmetry suggests architectural compatibility differences.
@@
| TreeMatching (medium) | 18.8M | 75% |
| TreeMatching (large) | 36M | 75.20% |
@@
TreeMatchingNet exhibits approximately the same performance between the medium model (18.8M parameters) and the large model (36M parameters), both achieving 75% accuracy despite a 2 increase in parameters. This plateau suggests an architectural bottleneck in the aggregation mechanism that prevents effective parameter utilization, regardless of the capacity added through additional propagation layers or hidden dimensions.
@@
The embedding model results illuminate why this was so challenging for BERT: BertEmbeddingNet achieves 45.78% without randomized pairing, while BertMatchingNet achieves only 35.38% with randomized pairing. This 10.4-point degradation indicates that BERT’s architecture cannot effectively learn from randomized pairs processed through cross-attention. The transformer’s attention mechanisms, optimized for learning from sequential structure, fail when forced to learn from mismatched pairs where the cross-attention target is deliberately chosen to be negative.
In contrast, TreeMatchingNet (75.20%) shows no degradation from TreeEmbeddingNet (57.57%), instead outperforming it by 17.63 points, indicating that graph-based message passing provides robustness to pairing randomization that transformers lack.
@@
The observed scaling plateau (18.8M params 36M params 75% accuracy) points to the aggregation mechanism as the primary source of information loss. The current pooling aggregation collapses node representations (each 1536-dimensional) into a single 2048-dimensional graph embedding through gated weighted sum pooling. This lossy compression represents the greatest potential for representation collapse in the architecture. While the GNN propagation layers learn rich, node-level distinctions about syntactic relationships and semantic content, the aggregation step treats all nodes similarly through a simple pooling operation, potentially destroying the structural information encoded during propagation.
@@
Both TMN models exhibit somewhat weaker performance on the Neutral class (65.00% recall) compared to Contradiction (81.19%) and Entailment (79.00%). This pattern reflects the inherent difficulty of the Neutral category.
@@ Conclusion
Key findings: (1) TreeMatchingNet achieves 75.20% SNLI accuracy vs BertMatchingNet’s 35.38%, demonstrating a clear performance advantage; (2) embedding model experiments show the structural advantage persists (TreeEmbedding 57.57% vs BertEmbedding 45.78%), with narrower gap indicating cross-attention amplifies structural benefits; (3) BertEmbeddingNet additionally outperforms BertMatchingNet, confirming that cross-attention interferes with BERT’s architecture; (4) increasing TMN parameters 2 (18.8M to 36M) yields minimal improvement, revealing a scaling bottleneck consistent with prior literature on structure-based NLP approaches, but which may be overcome by introducing scalable attention-based architectures into the aggregation step of TMN.
Implications: Explicit structural representations outperform sequence-based transformers at moderate scales, with benefits persisting across both matching and embedding paradigms. Inductive biases aligned with linguistic structure provide substantial learning efficiency gains. Cross-attention mechanisms amplify structural advantages for tree-based models but interfere with transformer architectures, particularly under randomized pairing regimes. Simple pooling aggregation creates bottlenecks preventing effective parameter scaling; attention-based aggregation may address this limitation.